OpenStudy (anonymous):

"The measure of the length of the top of a rectangular table is three times the measure of the width of the table. If the length was decreased by 3 feet and its width increased by 2 feet, the sides would be equal. What are the dimensions of the table top?" I'm pretty sure the starting equation thing is L=25+3w, but I could definitely be wrong.

OpenStudy (anonymous):

i) L = 3w ii) L - 3 = w + 2 i -> ii: 3w - 3 = w + 2 2w = 5 w = 2.5 i: L = 3w L = 3 * 2.5 L = 7.5